A rigid tank contains 61.5 g of chlorine gas (Cl2) at a temperature of 83 °C and an absolute pressure of 5.70 × 10^5 Pa. The mass per mole of Cl2 is 70.9 g/mol. Determine the volume of the tank.

Answer in liters please
Answer #1

Here,

let the volume of tank is V

P * V = n * R *

5.7 *10^5 * V = (61.5/70.9) * 8.314 * (83 + 273)

solving for V

V = 4.41 *10^-3 m^3 = 4.41 litres

the volume of tank is 4.41 litres